OpenStudy (anonymous):

Let f(x) = x^2 + 6 and g(x) =x+8/x. Find ( g o f)(-7)

OpenStudy (agent0smith):

find f(-7) first, ie plug in x=-7 into f(x) then plug that result into g(x)

OpenStudy (anonymous):

f(x)=-7^2+6?

OpenStudy (agent0smith):

Yes. Brackets around -7.

OpenStudy (anonymous):

55?

OpenStudy (agent0smith):

Yes, now plug that into g(x)

OpenStudy (anonymous):

x would be 55 in g(x) ?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

55+8/55? i got 1.14

OpenStudy (agent0smith):

I have no idea, you need to make sure to use brackets for g(x) =x+8/x because I can't tell what it is like that

OpenStudy (anonymous):

g(x)= (55)+8/(55)? So i would get 63/55?

OpenStudy (anonymous):

Thats my answer right? :)

OpenStudy (agent0smith):

No, i mean, make it clear what it is. is g(x) =x + 8/x or g(x) =(x+8)/x

OpenStudy (agent0smith):

\[\large g(x) =x+\frac{ 8 }{x} \text { ..... or ..... } g(x) =\frac{x+8 }{x} \]

OpenStudy (anonymous):

second one

OpenStudy (agent0smith):

Then use brackets, g(x) =(x+8)/x, since the g(x) =x + 8/x means the first one.

OpenStudy (anonymous):

So 63/55 isnt my answerr?

OpenStudy (anonymous):

lol its one of my choices

OpenStudy (agent0smith):

Plug in 55 into your function, the correct one, and check :P

OpenStudy (anonymous):

g(x)= (55+8)/55 ?

OpenStudy (anonymous):

Its the same thing right?

OpenStudy (agent0smith):

Yes, simplify it

OpenStudy (anonymous):

But thats my answer like that just like that 63/55 :P Its A:)

OpenStudy (agent0smith):

lol yes, i was just making sure you KNEW it was your answer! :P

OpenStudy (anonymous):

oh haha okay :) thanks :)

OpenStudy (agent0smith):

haha no prob :)